OpenStudy (griffindore):

Mindy conducted a survey of the phones in four stores and wrote the following observations: 5 of the 8 phones in Store 1 were smartphones 13 of the 22 phones in Store 2 were smartphones 22 of the 30 phones in Store 3 were smartphones 12 of the 25 phones in Store 4 were smartphones Based on the survey, which store had the greatest percentage of smart phones?

OpenStudy (anonymous):

1.5/8=62.5% 2.13/22=59% 3.22/30=73% 4.12/25=48% Which one has the greatest percentage of smartphones?

OpenStudy (griffindore):

i might know

OpenStudy (griffindore):

hold on

OpenStudy (griffindore):

oh sorry never mind

OpenStudy (anonymous):

okay which is the greatest or biggest number 62.5, 59, 73 and 48

OpenStudy (anonymous):

do you know?

OpenStudy (anonymous):

hello???